OpenStudy (anonymous):

Factor 8x2 + 12x – 3. A. (2x – 1)(4x + 3) B. (2x + 3)(4x – 1) C. (8x + 1)(x – 3) D. The polynomial is prime.

OpenStudy (anonymous):

You cannot factor it ove the integers or rationals

OpenStudy (anonymous):

It should be prime

OpenStudy (anonymous):

It has real roots, but they are irrationals

OpenStudy (anonymous):

The roots are \[ \left\{\left\{x= \frac{1}{4} \left(-3-\sqrt{15}\right)\right\},\left\{x= \frac{1}{4} \left(\sqrt{15}-3\right)\right\}\right\} \]